OpenStudy (anonymous):

few questions pretty smiple how do I multiply x^2 times X^2 and how do I multiply x^2 times 3

Mehek (mehek14):

\(x^2*x^2\\ x^2*x^3\) like that?

Mehek (mehek14):

@rebeccahumbarger ?

OpenStudy (anonymous):

yes

Mehek (mehek14):

to multiply \(x^2*x^2\) all you have to do is add the exponents 2 + 2 = 4 so \(x^2*x^2=x^4\) do you understand? @rebeccahumbarger ?

Mehek (mehek14):

you would do the same thing with \(x^2*x^3\) add 2 + 3 and keep the same base